class contains 16 boys and 10 girls of which half of the girls are fat. Find the probability that a student chosen at random is a boy or a fat student.



1
Expert's answer
2021-07-27T16:16:44-0400
"P(boy\cup fat)=\dfrac{16+\dfrac{1}{2}(10)}{16+10}=\dfrac{21}{26}"